OpenStudy (anonymous):

Determine the convergence or divergence of the p-series $\displaystyle\sum\limits_(n=1)^Infinity 1/n^(ln3) showing all work. Can someone help me out? And I'm just learning latex, so bear with me.

OpenStudy (anonymous):

\pi n=1 to infinity of 1/n^(ln3)

OpenStudy (anonymous):

ln 3 > 1

OpenStudy (anonymous):

So it is convergent

OpenStudy (anonymous):

By the p-series test

OpenStudy (anonymous):

\[\sum_{1}^{\infty}\frac{1}{n^{\ln3}}\]\[\ln3 \approx1.0986\]A p series convergeses if p > 1. As shown above, ln3 is greater than 1.

OpenStudy (anonymous):

Great, thank you both!!